OpenStudy (anonymous):

A 70 ft ladder is mounted 10 ft above the ground on a fire truck. The bottom of the ladder is 40 ft. From the wall of a building. The top of the ladder is touching the building. How high off the ground is the top of the ladder?

OpenStudy (anonymous):

Right triangle with hypotenuse. = 70 + 10 = 80 One side: 40. The second side is x. Use the pythagor formula: x^2 + (40)^2 = (80)^2 --> x = ...